# Towards Enhancing Time Series Contrastive Learning: A Dynamic Bad Pair Mining Approach

### v2

@@ 1 Introduction
Self-supervised contrastive learning has shown remarkable efficacy in learning meaningful representations from unlabeled time series data, thereby improving the performance of downstream tasks ($e.g.$, time series classification) (Oord et al., 2018; Tonekaboni et al., 2021; Eldele et al., 2021; Yèche et al., 2021; Lan et al., 2022; Yue et al., 2022; Yang & Hong, 2022; Zhang et al., 2022; Ozyurt et al., 2023). The key concept of time series contrastive learning is to capture the meaningful underlying information shared between views (usually generated by random data augmentation), such that the learned representations are discriminative among time series. To accomplish this objective, existing contrastive methods comply with an assumption that the augmented views from the same instance ($i.e.$, positive pair) share meaningful semantic information (Chen et al., 2020; He et al., 2020a; Chen & He, 2021a; Wen & Li, 2021).
@@ 2.2 Time Series Contrastive Learning
Several recent studies have demonstrated that contrastive learning is a prominent self-supervised approach for time series representation learning. For example, TSTCC (Eldele et al., 2021) proposed temporal and contextual contrasting that worked with a weak-strong augmentation strategy to learn discriminative time series representation. BTSF (Yang & Hong, 2022) presented a bilinear temporal-spectral fusion module that utilized the temporal-spectral affinities to improve the expressiveness of the representations. CoST (Woo et al., 2022) applied contrastive learning to learn disentangled seasonal-trend representations for long sequence time series forecasting. TS2Vec (Yue et al., 2022) utilized hierarchical contrasting over augmented context views such that the learned contextual representation for each timestamp is robust. TF-C (Zhang et al., 2022) leveraged time-frequency consistency as a mechanism in the pre-training to facilitate knowledge transfer between datasets.
@@ 3.2 Analysis: Time Series Contrastive Learning
Noisy Alignment. Let us consider the noisy positive pair problem. According to (Wen & Li, 2021), each input data can be represent in the form of $x_{i}=z_{i}+\xi_{i}$, where $z_{i}\sim\mathcal{D}_{z}$ denotes the true signal that contains the desired features we want the encoder $G(\theta)$ to learn, and $\xi_{i}\sim\mathcal{D}_{\xi}$ is the spurious dense noise. We hypothesize that the noisy positive pair problem is likely to occur on $x_{i}$ when $\xi_{i}\gg z_{i}$, because noisy raw signals will produce noisy views after data augmentation. Formally, we define $(u_{i},v_{i})$ as a noisy positive pair when $\xi_{i}^{u}\gg z_{i}^{u}$ and $\xi_{i}^{v}\gg z_{i}^{v}$. Further, we know that minimizing the contrastive loss is equivalent to maximize the mutual information between the representations of two views (Tian et al., 2020). Therefore, when noisy positive pairs present, $\operatorname*{arg\,max}_{\theta}(I(G(u_{i}|\theta);G(v_{i}|\theta)))$ is approximate to $\operatorname*{arg\,max}_{\theta}(I(G(\xi_{i}^{u}|\theta);G(\xi_{i}^{v}|\theta)))$. This results in noisy alignment, where the model predominantly learning patterns from noise.
@@ 4.1 Datasets, Tasks, and Baselines
Our model is evaluated on four real-world benchmark time series datasets: PTB-XL (Wagner et al., 2020) is to-date largest freely accessible clinical 12-lead ECG-waveform dataset. Based on ECG annotation method, there are three multi-label classification tasks: Diagnostic (44 classes), Form (19 classes), and Rhythm (12 classes). HAR (Anguita et al., 2013): multi-class classification for 6 different activities. Sleep-EDF (Goldberger et al., 2000): multi-class classification for 5 sleep stages. Epilepsy (Andrzejak et al., 2001): binary classification to identify epileptic seizure. We use AUROC as test metric for PTB-XL as in (Strodthoff et al., 2020), Accuracy and AUPRC for HAR, Sleep-EDF, and Epilepsy as in (Eldele et al., 2021). See Appendix A.2 for detailed descriptions of each dataset.
@@ 4.2 Linear Evaluation
 + DBPM 81.56±${}_{\pm 1.84}$ 78.07±${}_{\pm 1.01}$ 80.89±${}_{\pm 3.85}$
extra run-time ${}_{(+0.74s/epoch)}$ ${}_{(+0.42s/epoch)}$ ${}_{(+0.54s/epoch)}$
@@
 + RINCE $72.85_{\pm 1.79}$ 68.49±${}_{\pm 1.82}$ $74.98_{\pm 3.70}$
 + DBPM 77.32±${}_{\pm 1.65}$ $68.28_{\pm 1.16}$ 80.31±${}_{\pm 0.47}$
extra run-time ${}_{(+0.40s/epoch)}$ ${}_{(+0.37s/epoch)}$ ${}_{(+0.89s/epoch)}$
@@
 + DBPM 67.00±${}_{\pm 1.42}$ 63.90±${}_{\pm 1.34}$ 69.14±${}_{\pm 1.59}$
extra run-time ${}_{(+0.92s/epoch)}$ ${}_{(+0.64s/epoch)}$ ${}_{(+0.93s/epoch)}$
@@
 + DBPM 77.46±${}_{\pm 0.37}$ 69.81±${}_{\pm 0.42}$ 84.70±${}_{\pm 0.17}$
extra run-time ${}_{(+0.27s/epoch)}$ ${}_{(+0.31s/epoch)}$ ${}_{(+0.73s/epoch)}$
@@
 + DBPM 74.04±${}_{\pm 1.03}$ 72.03±${}_{\pm 1.92}$ 82.10±${}_{\pm 1.88}$
extra run-time ${}_{(+0.94s/epoch)}$ ${}_{(+0.39s/epoch)}$ ${}_{(+1.52s/epoch)}$
@@
 + RINCE $69.37_{\pm 2.31}$ $70.02_{\pm 2.28}$ 68.22±${}_{\pm 3.41}$
 + DBPM 73.32±${}_{\pm 3.08}$ 71.67±${}_{\pm 0.71}$ $67.90_{\pm 3.66}$
extra run-time ${}_{(+1.34s/epoch)}$ ${}_{(+0.26s/epoch)}$ ${}_{(+0.81s/epoch)}$
[…]
@@
 + DBPM 85.70±${}_{\pm 0.74}$ 92.97±${}_{\pm 0.56}$ 74.37±${}_{\pm 0.21}$ 69.10±${}_{\pm 0.37}$ 95.85±${}_{\pm 0.14}$ 98.15±${}_{\pm 0.10}$
[…]
@@ Appendix A.2 Dataset Descriptions
HAR. In the UCI Human Activity Recognition dataset (Anguita et al., 2013), data is collected from wearable sensors installed on 30 subjects participating in 6 different activities, where each recording contains 9 time series with length of 128. Epilepsy. The Epileptic Seizure Recognition dataset (Andrzejak et al., 2001) consists of brain activity recordings from 500 subjects, where each recording contains single time series with length of 178. Sleep-EDF (Goldberger et al., 2000) contains whole-night Polysomnography (PSG) sleep recordings categorized into 5 sleep stages. Each recording contains single time series with length of 3,000. We follow the data pre-processing steps from (Eldele et al., 2021) for these three datasets. Table 4 provides a summarization of HAR, Epilepsy, and Sleep-EDF.
@@ Appendix A.4 Gradient Derivation of InfoNCE Loss
To simplify the notation, let us denote $\operatorname{exp}(\bm{r}_{i}^{u\top}\bm{r}_{0}^{v-}/t)=\operatorname{exp}(\bm{r}_{i}^{u\top}\bm{r}_{i}^{v}/t)$. Therefore, Eq. 12 can be expressed as:

### v3

@@ 1 Introduction
Self-supervised contrastive learning has shown remarkable efficacy in learning meaningful representations from unlabeled time series data, thereby improving the performance of downstream tasks ($e.g.$, time series classification) (Oord et al., 2018; Tonekaboni et al., 2021; Eldele et al., 2021; Yèche et al., 2021; Lan et al., 2022; Yue et al., 2022; Yang & Hong, 2022; Zhang et al., 2022b; Ozyurt et al., 2023). The key concept of time series contrastive learning is to capture the meaningful underlying information shared between views (usually generated by random data augmentation), such that the learned representations are discriminative among time series. To accomplish this objective, existing contrastive methods comply with an assumption that the augmented views from the same instance ($i.e.$, positive pair) share meaningful semantic information (Chen et al., 2020; He et al., 2020a; Chen & He, 2021a; Wen & Li, 2021).
@@ 2.2 Time Series Contrastive Learning
Several recent studies have demonstrated that contrastive learning is a prominent self-supervised approach for time series representation learning. For example, TSTCC (Eldele et al., 2021) proposed temporal and contextual contrasting that worked with a weak-strong augmentation strategy to learn discriminative time series representation. BTSF (Yang & Hong, 2022) presented a bilinear temporal-spectral fusion module that utilized the temporal-spectral affinities to improve the expressiveness of the representations. CoST (Woo et al., 2022) applied contrastive learning to learn disentangled seasonal-trend representations for long sequence time series forecasting. TS2Vec (Yue et al., 2022) utilized hierarchical contrasting over augmented context views such that the learned contextual representation for each timestamp is robust. TF-C (Zhang et al., 2022b) leveraged time-frequency consistency as a mechanism in the pre-training to facilitate knowledge transfer between datasets.
@@ 3.2 Analysis: Time Series Contrastive Learning
Noisy Alignment. Let us consider the noisy positive pair problem. According to Wen & Li (2021), each input data can be represent in the form of $x_{i}=z_{i}+\xi_{i}$, where $z_{i}\sim\mathcal{D}_{z}$ denotes the true signal that contains the desired features we want the encoder $G(\theta)$ to learn, and $\xi_{i}\sim\mathcal{D}_{\xi}$ is the spurious dense noise. We hypothesize that the noisy positive pair problem is likely to occur on $x_{i}$ when $\xi_{i}\gg z_{i}$, because noisy raw signals will produce noisy views after data augmentation. Formally, we define $(u_{i},v_{i})$ as a noisy positive pair when $\xi_{i}^{u}\gg z_{i}^{u}$ and $\xi_{i}^{v}\gg z_{i}^{v}$. Further, we know that minimizing the contrastive loss is equivalent to maximize the mutual information between the representations of two views (Tian et al., 2020). Therefore, when noisy positive pairs present, $\operatorname*{arg\,max}_{\theta}(I(G(u_{i}|\theta);G(v_{i}|\theta)))$ is approximate to $\operatorname*{arg\,max}_{\theta}(I(G(\xi_{i}^{u}|\theta);G(\xi_{i}^{v}|\theta)))$. This results in noisy alignment, where the model predominantly learning patterns from noise.
@@ 4.1 Datasets, Tasks, and Baselines
Our model is evaluated on four real-world benchmark time series datasets: PTB-XL (Wagner et al., 2020) is to-date largest freely accessible clinical 12-lead ECG-waveform dataset. Based on ECG annotation method, there are three multi-label classification tasks: Diagnostic (44 classes), Form (19 classes), and Rhythm (12 classes). HAR (Anguita et al., 2013): multi-class classification for 6 different activities. Sleep-EDF (Goldberger et al., 2000): multi-class classification for 5 sleep stages. Epilepsy (Andrzejak et al., 2001): binary classification to identify epileptic seizure. We use AUROC as test metric for PTB-XL as in Strodthoff et al. (2020), Accuracy and AUPRC for HAR, Sleep-EDF, and Epilepsy as in Eldele et al. (2021). See Appendix A.2 for detailed descriptions of each dataset.
@@ 4.2 Linear Evaluation
   + DBPM 81.56±1.84 78.07±1.01 80.89±3.85
extra run-time (+0.74s/epoch) (+0.42s/epoch) (+0.54s/epoch)
@@
   + RINCE $72.85_{\pm 1.79}$ 68.49±1.82 $74.98_{\pm 3.70}$
   + DBPM 77.32±1.65 $68.28_{\pm 1.16}$ 80.31±0.47
extra run-time (+0.40s/epoch) (+0.37s/epoch) (+0.89s/epoch)
@@
   + DBPM 67.00±1.42 63.90±1.34 69.14±1.59
extra run-time (+0.92s/epoch) (+0.64s/epoch) (+0.93s/epoch)
@@
   + DBPM 77.46±0.37 69.81±0.42 84.70±0.17
extra run-time (+0.27s/epoch) (+0.31s/epoch) (+0.73s/epoch)
@@
   + DBPM 74.04±1.03 72.03±1.92 82.10±1.88
extra run-time (+0.94s/epoch) (+0.39s/epoch) (+1.52s/epoch)
@@
   + RINCE $69.37_{\pm 2.31}$ $70.02_{\pm 2.28}$ 68.22±3.41
   + DBPM 73.32±3.08 71.67±0.71 $67.90_{\pm 3.66}$
extra run-time (+1.34s/epoch) (+0.26s/epoch) (+0.81s/epoch)
extra memory ${}_{+1.40\textit{MB}}$ ${}_{+0.57\textit{MB}}$ ${}_{+1.40\textit{MB}}$
@@
   + DBPM 85.70±0.74 92.97±0.56 74.37±0.21 69.10±0.37 95.85±0.14 98.15±0.10
extra memory ${}_{+0.58\textit{MB}}$ ${}_{+2.60\textit{MB}}$ ${}_{+0.72\textit{MB}}$
@@ Appendix A.2 Dataset Descriptions
HAR. In the UCI Human Activity Recognition dataset (Anguita et al., 2013), data is collected from wearable sensors installed on 30 subjects participating in 6 different activities, where each recording contains 9 time series with length of 128. Epilepsy. The Epileptic Seizure Recognition dataset (Andrzejak et al., 2001) consists of brain activity recordings from 500 subjects, where each recording contains single time series with length of 178. Sleep-EDF (Goldberger et al., 2000) contains whole-night Polysomnography (PSG) sleep recordings categorized into 5 sleep stages. Each recording contains single time series with length of 3,000. We follow the data pre-processing steps from Eldele et al. (2021) for these three datasets. Table 4 provides a summarization of HAR, Epilepsy, and Sleep-EDF.
@@ Appendix A.4 Gradient Derivation of InfoNCE Loss
To simplify the notation (Zhang et al., 2022a), let us denote $\operatorname{exp}(\bm{r}_{i}^{u\top}\bm{r}_{0}^{v-}/t)=\operatorname{exp}(\bm{r}_{i}^{u\top}\bm{r}_{i}^{v}/t)$. Therefore, Eq. 12 can be expressed as: